show work please

What is the compound annual interest rate that turns a $1,000 initial capital into a $5200
cumulative investment after 12 years?
Answer #1

F = P *(1+i)^t

here F = 5200

P = 1000

t = 12

putting values into equation

5200 = 1000 * (1+i)^12

(1+i)^12 = 5200/1000 = 5.2

(1+i) = 5.2^(1/12)

1+i = 1.147273

i = 0.147273 = 14.73%